At what temperature in OC, will a 2.42msolution of ethanol in water freeze? Kfpwater=1.858OC/m

Do NOT write units with your answer!!!

If 11.57 of an unknown covalent compound is placed in 100.0mL of water. The solution boils at 101.182oC, what is the molar mass of the covalent compound in g/mol ?

Kbpwater=0.512OC/m   Density of water is 1.00g/mL.

Do NOT write units with your answer!!!

If 7.29 of an unknown covalent compound is placed in 100.0mL of water. The solution boils at 100.205oC, what is the molar mass of the covalent compound in g/mol ?

Kbpwater=0.512OC/m   Density of water is 1.00g/mL.

Do NOT write units with your answer!!!

Choose ALL Correct:

A colligative properties depend on ...

The concentration of solute particles.

The identity of the solute.

The identity of the solvent.

Chemical reactions between the solute and solvent.
Answer #1

i)

Given data, Kf=1.858 C/mol

m=2.42

Let's Consider

\DeltaTf = Kf*m

= 1.858*2.42

  \DeltaTf = 4.496 degree C

ii

boiling point of water = 100 oC

boiling point of mixture = 101.182oC

dTb = 101.182 - 100

=1.18 oC

mass of unknown covalent compound = 11.57 g

density of water = 1 g/ml

mass of water = 100 ml x 1 g/ml

= 100 g

= 0.100 kg

molality of solution (m) = 11.57g/molar mass x 0.100 kg

using, dTb = iKbm

i = 1 (Van't hoff factor)

Kb for water = 0.512 oC/m

Substitute these value in above formula,

1.18=1x0.512 x 11.57/molar mass x 0.100

Molar mass of an covalent compound is 50.2 g/mol

iii)

boiling point of water = 100 oC

boiling point of mixture = 100.205oC

dTb = 100.205oC - 100

= 0.205 oC

mass of unknown covalent compound = 7.29g

density of water = 1 g/ml

mass of water = 100 ml x 1 g/ml

= 100 g

= 0.100 kg

molality of solution (m) = 7.29g/molar mass x 0.100 kg

i = 1 (Van't hoff factor)

Kb for water = 0.512 oC/m

using,

dTb = iKbm

0.205 = 1 x 0.512 x 7.29/molar mass x 0.100

Molar mass of a covalent compound is 1.82 g/mol

iv)

A colligative properties depend on The concentration of solute particles

option 1 is correct